OpenStudy (shadyboy):

3x = 36 A x=9 B x=12 C x=16 D x=18

OpenStudy (igreen):

Divide 3 to both sides. 36 / 3 = ?

OpenStudy (haseeb96):

step #1 3x=36 step # 2 x =36/3 x= 12